OpenStudy (anonymous):

Calc 2 Polar Equations Question: What is the area from 0 to pi/2 of r = sin(2θ)

OpenStudy (anonymous):

\[\Large Area=\int\limits_{0}^{\frac{\pi}{2}}\sqrt{r^2 + (\frac{dr}{d \theta})^2}d \theta \] \[\Large r = sin(2 \theta )\] \[\Large (\frac{dr}{d \theta})^2 = 2cos(2 \theta)\] So \[\Large Area=\int\limits_{0}^{\frac{\pi}{2}}\sqrt{sin^2(2 \theta ) + (2cos(2 \theta))^2}d \theta \] And at this point I can't figure out how to manipulate the equation into a workable integral... :/

OpenStudy (anonymous):

That's not the area integral. It should be \[A = \int_a^b \frac{1}{2} r(\theta)^2 d\theta \]

OpenStudy (anonymous):

What you have there is the arc length integral.

OpenStudy (anonymous):

doh. >.< That would probably make things a lot easier. Thanks.